# Domain Transfer Through Image-to-Image Translation for Uncertainty-Aware Prostate Cancer Classification

### v1

@@ Abstract
Prostate Cancer (PCa) is often diagnosed using High-resolution 3.0 Tesla(T) MRI, which has been widely established in clinics. However, there are still many medical centers that use 1.5T MRI units in the actual diagnostic process of PCa. In the past few years, deep learning-based models have been proven to be efficient on the PCa classification task and can be successfully used to support radiologists during the diagnostic process. However, training such models often requires a vast amount of data, and sometimes it is unobtainable in practice. Additionally, multi-source MRIs can pose challenges due to cross-domain distribution differences. In this paper, we have presented a novel approach for unpaired image-to-image translation of prostate mp-MRI for classifying clinically significant PCa, to be applied in data-constrained settings. First, we introduce domain transfer, a novel pipeline to translate unpaired 3.0T multi-parametric prostate MRIs to 1.5T, to increase the number of training data. Second, we estimate the uncertainty of our models through an evidential deep learning approach; and leverage the dataset filtering technique during the training process. Furthermore, we introduce a simple, yet efficient Evidential Focal Loss that incorporates the focal loss with evidential uncertainty to train our model. Our experiments demonstrate that the proposed method significantly improves the Area Under ROC Curve (AUC) by over 20% compared to the previous work (98.4% vs. 76.2%). We envision that providing prediction uncertainty to radiologists may help them focus more on uncertain cases and thus expedite the diagnostic process effectively. Our code is available at https://github.com/med-i-lab/DT_UE_PCa.
Keywords: Deep Learning, Image Translation, Uncertainty Estimation, Prostate Cancer
@@ 1 Introduction
Prostate Cancer (PCa) is a prevalent form of cancer among men (Reda et al., 2018), and the clinically significant PCa is defined by the Gleason score $>$ 6 or the histopathology ISUP grade $\geq$ 2 (Smith et al., 2004; Arif et al., 2020). The current PCa diagnosis procedure involves a combination of the prostate-specific antigen test and the histopathology analysis of the Transrectal Ultrasound-guided biopsy (TRUS) taken from 10-12 regions on the prostate gland (Fletcher, 2019; Grebenisan et al., 2021). However, the histopathology analysis on TRUS can miss up to 20% of clinically significant PCa due to the limited number of biopsy samples (Grebenisan et al., 2021; Reda et al., 2018). Multi-parametric Magnetic Resonance Imaging (mp-MRI) has emerged as an effective alternative to TRUS for the early detection of PCa. mp-MRI uses a combination of anatomical and functional sequences of MRI that can further highlight the differences between normal and abnormal (cancer) cells. 3.0T MRI generally has higher image quality and spatial resolution (Ladd et al., 2018) than 1.5T MRI, but the latter is widely used in local, small clinical centers due to its lower price (Mushlin et al., 1997). The evaluation and reporting guideline of prostate mp-MRI was first introduced in the Prostate Imaging Reporting and Data System (PI-RADS) (Barentsz et al., 2012; Weinreb et al., 2016; Barentsz et al., 2016). The guideline provides a comprehensive scoring schema for suspicious prostate lesions and mp-MRI sequences. An extensive Prostate MRI imaging study (PROMIS) (Bosaily et al., 2015) reported that targeted biopsy using mp-MRI has higher sensitivity and negative-predictive value (NPV) but lower specificity compared to TRUS biopsy (Bosaily et al., 2015; Stabile et al., 2020; Ahmed et al., 2017). The study also showed that 27% of the patients did not need to undergo biopsy, had mp-MRI been used for screening. Although PROMIS provides strong practical implications for mp-MRI in PCa diagnosis, the low specificity indicates that mp-MRI can be plausibly improved by advanced analyses.
@@
In recent years, deep learning methods have emerged as a powerful tool for image classification tasks, and have provided promising performance in detecting and segmenting PCa on multi-parametric Prostate MRIs (Saha et al., 2021; Le et al., 2017; Yoo et al., 2019; Iqbal et al., 2021; Pellicer-Valero et al., 2022). A more recent grand challenge, ProstateX (Armato et al., 2018), has further shown the ability of deep learning approaches in detecting clinically significant PCa on 3.0T mp-MRI data. Several groups have developed Convolutional Neural Network (CNN)-based models that achieve high performance for PCa classification (Litjens et al., 2014; Liu et al., 2017; Mehrtash et al., 2017; Armato et al., 2018; Grebenisan et al., 2020, 2021). These methods have a great potential for clinical translations by highlighting abnormal lesions for radiologists during the PCa diagnostic process.
While deep learning has shown promising results in detecting PCa on mp-MRI, there are several challenges in deploying deep models in local clinics with limited data and patient throughput. Training deep models typically requires a large amount of data, making them difficult to deploy in small clinics. Moreover, MRI data may be acquired under different magnetic strengths, vendors, and protocols, which can affect the performance of deep models. For example, prostate MRI typically with high magnetic strength of 3.0T (Ullrich et al., 2017) is preferred because it produces high-resolution images and provides detailed information. In contrast, the low magnetic strength MRI (1.5T) may result in fuzzy boundaries (Ladd et al., 2018) and not be able to offer detailed information. The performance of deep models will be significantly affected if there is a difference between training and test distribution. There are efforts in the literature on solving related problems in federated learning (Li et al., 2020; Adnan et al., 2022) which is not a focus on our study. Furthermore, classical deep models are designed to predict a label when inferring data from a test set, regardless of whether or not the test image is in or out of the training set distribution. These models are not able to identify data samples that belong to unrelated distributions (Sensoy et al., 2018), or indicate how confident they are in their prediction. These limitations make models hard to interpret and hence, there are concerns about the reliability of such models. Hence, reusing and deploying models for local PCa detection is challenging. It is essential to address the above limitations and drawbacks when deploying deep learning models to real clinical routines. Thus, two main questions arise in this context:
@@
    For small local clinical centers, can they take advantage of the large high-resolution 3.0T public MRI data and enhance the classification performance on their limited low-resolution local 1.5T MRI data?
@@
    When deploying models in clinical centers, could we offer additional information regarding the confidence of the model’s predictions, in addition to the final result, to enhance the reliability of the models?
In this work, we aim to answer the two questions above. We propose a novel 2-stage learning framework for clinically significant PCa classification using multi-parametric, multi-center MRI data that can simultaneously provide an estimate of the predictive confidence and the corresponding predicted label to improve classification performance. In the first stage, we introduce a data preprocessing pipeline that translates prostate mp-MRI data from 3.0T to 1.5T via a Generative Adversarial Network (GAN) approach in order to increase the number of training samples. This step addresses the challenge of limited data in local clinics with low patient throughput (see Section 4.1). In the second stage, we propose an uncertainty-aware PCa classification approach. Specifically, we design three different model architectures and leverage the co-teaching framework (Han et al., 2018) to address the noisy label problem (see Section 4.2.1). During the training phase, we incorporate dataset filtering using evidential uncertainty estimation (Sensoy et al., 2018) to eliminate the training data samples with high prediction uncertainty to improve the robustness of our models. Finally, we extend the work of Sensoy et al. (2018) to design a novel Evidential Focal Loss to optimize our classification models during training (see Section 4.2.2). Experiments demonstrate the effectiveness of the proposed framework in significantly improving the classification performance compared to previous work.
@@
    We develop a GAN-based framework to translate unpaired prostate mp-MRIs from 3.0T to 1.5T, which we termed as domain transfer. This framework would align different data distributions and increase the number of training data for deep classification models.
@@
    We incorporate the Theory of Evidence (Yager and Liu, 2008) into our model, enabling it to identify and filter out highly uncertain training data and making the model more robust. We propose a novel loss function termed Evidential Focal Loss that combines the original Focal Loss (Lin et al., 2017) and the evidential uncertainty (Sensoy et al., 2018) for the binary PCa classification task.
@@
    Using the uncertainty and filtering on the training set, our results outperform the state-of-the-art and improve the interpretability of model predictions. By providing confidence estimates for the predictions, radiologists can make informed decisions during the PCa diagnostic process and effectively expedite the process.
@@ 2.1 Domain Adaptation
Machine learning algorithms usually perform well when training and test data share the same distribution and feature space. However, in real-world applications, the distribution of test data often shifts, leading to biased or inaccurate predictions. In addition, it is time-consuming or infeasible to acquire new training data and fully repeat training steps. Domain Adaptation (DA) is an approach that addresses this issue by mitigating the dataset bias or domain shift problem caused by different distributions. There has been a lot of work on this topic in the past few years, which can be grouped into the following three general tasks (Cui et al., 2020): (1) unsupervised DA tasks (Ganin and Lempitsky, 2015; Ganin et al., 2016; Long et al., 2016; Saito et al., 2018; Long et al., 2014) focus on addressing the domain shift problem without requiring labeled target domain data; (2) semi-supervised DA tasks(Yao et al., 2015; Saito et al., 2019; Li et al., 2021) aim to explore the partially labeled target domain data to further enhance the performance of domain adaptation algorithms; and (3) multi-source DA tasks (Hoffman et al., 2012; Xu et al., 2018; Peng et al., 2019) deal with scenarios where multiple source domains are available for adaptation. DA methods are often used to extract domain-invariant features for transferring knowledge between source and target domains. These methods incorporate various learning objectives with deep neural networks Wang and Deng (2018) for distribution matching: (1). Discrepancy Measurement-based methods aim to align feature distributions between two domains by fine-tuning deep models, e.g., using statistic criterion like Maximum Mean Discrepancy (Long et al., 2015; Yan et al., 2017; Kumagai and Iwata, 2019), and class criterion (Tzeng et al., 2015; Hinton et al., 2015; Motiian et al., 2017). Some of these methods often require large labeled target domain data to diminish the domain shift problem, which is sometimes infeasible to get such medical data in the real-life scenario. (2). Adversarial-based methods aim to confuse domain discriminators from Generative Adversarial Networks (GANs) to enhance the invariant feature extraction (Ganin et al., 2016; Bousmalis et al., 2017; Hong et al., 2018). One common scenario involves utilizing noise vectors, either with or without source images, to generate realistic target images while preserving the source features. However, training GANs are hard and sometimes results in generator degradation, e.g., mode collapse (Karras et al., 2020). (3). Reconstruction-based methods, in addition to the general GANs approach from the above category, aim to reconstruct source-like images as an auxiliary task to preserve domain invariant features through an adversarial reconstruction paradigm (Hoffman et al., 2018; Zhu et al., 2017). These methods usually have superior performance over the conventional GANs approach because they have an explicit reconstruction task to supervise the entire pipeline and make the training process more stable.
@@
CycleGAN (Zhu et al., 2017) is one of the state-of-the-art unsupervised adversarial reconstruction-based methods that is widely used for unpaired image-to-image translation. Its cycle consistency loss ensures the pixel-level similarity between two images through a reconstruction task, i.e., the source image $s$ is translated to the target domain $\hat{s}$ and then translated back $\tilde{s}$, where it should be identical to the original image ($s=\tilde{s}$). However, one drawback of cycle consistency loss is the harsh constraint on the pixel-level, which will degrade the performance of GANs in some tasks (Zhao et al., 2020). To address this limitation, Zhao et al. (2020) purpose the adversarial consistency loss GAN (ACL-GAN) that replaces the pixel-level similarity with the distance between distributions, i.e., instead of forcing $s=\tilde{s}$, we let the distribution of $\tilde{s}$ to be similar to the distribution of $s$. The ACL-GAN can retain important features from the source images and overcomes the disadvantage of the cycle-consistency constraint. Therefore, we adapt the ACL-GAN model and build our framework based on it.
@@
In medical imaging, domain shift problems usually fall into two variations: subject-related variation (age, gender, etc.), and acquisition-related variation (MRI vendor, field strength, imaging protocol, etc.) (Kouw et al., 2017). To solve the problem, one intuitive approach is to fine-tune a model that is pre-trained on the source domain with the new data from the target domain. Khan et al. (2019) propose to use the pre-trained VGG model on the ImageNet dataset (Deng et al., 2009) to learn robust high-level features of natural images, and then fine-tune it on the labeled MR images for the Alzheimer’s Disease (AD) classification task to achieve state-of-the-art performance. Similarly, Ghafoorian et al. (2017) study the impact of the fine-tuning techniques on the brain lesion segmentation task, demonstrating that fine-tuning with only a small number of target domain training samples can outperform models trained from scratch. Another approach is to use domain adaptation as an intermediate step to reduce variance in image acquisition parameters from both domains and then use it for downstream tasks. Researchers have attempted to address the problem of acquisition variation in MRI data for several years. Kouw et al. (2017) propose a feature-level representation learning method to either extract acquisition-invariant features or remove acquisition-variant features from paired 1.5T and 3.0T brain MRIs. The learned features are then used for a downstream classification task. However, obtaining paired 1.5T and 3.0T MRI data in real-life scenarios is impractical. Another way to align acquisition-invariant features is to synthesize images from different types of acquisition parameters using GAN-based adversarial reconstruction methods. GANs have been applied to perform cross-modality image translation between different medical images or generate synthetic images from random noise. The objective of such translation tasks is to retain the underlying structure while changing the appearance of the image (Armanious et al., 2020). Researchers have attempted to estimate images in the target modality from the source modality, such as MRI-CT translation (Hiasa et al., 2018; Nie et al., 2017; Oulbacha and Kadoury, 2020; Armanious et al., 2019), and X-ray to CT translation (Ying et al., 2019; Ge et al., 2022). Other areas that have been explored include intra-modality translation, such as 3.0T-7.0T MRI translation (Nie et al., 2018), T1/T2-FLAIR translation (Hu, 2021; Uzunova et al., 2020) and pure data augmentation by generating synthetic images from random noise vectors (Radford et al., 2015; Frid-Adar et al., 2018; Huang and Jafari, 2021; Kwon et al., 2019). However, most of the works do not consider the real clinical practicality, for example, for 3.0T-7.0T MRI translation in Nie et al. (2018), the training data is paired, which is not feasible in real clinical settings. Generating synthetic images from noise does not take advantage of the publicly available data and ignores a-priori information. The current limitations provide great potential for unpaired image translation for medical images, which we employ in this work.

### v2

@@ Abstract
Purpose: Prostate Cancer (PCa) is a prevalent disease among men, and multi-parametric MRIs offer a non-invasive method for its detection. While MRI-based deep learning solutions have shown promise in supporting PCa diagnosis, acquiring sufficient training data, particularly in local clinics remains challenging. One potential solution is to take advantage of publicly available datasets to pre-train deep models and fine-tune them on the local data, but multi-source MRIs can pose challenges due to cross-domain distribution differences. These limitations hinder the adoption of explainable and reliable deep-learning solutions in local clinics for PCa diagnosis. In this work, we present a novel approach for unpaired image-to-image translation of prostate multi-parametric MRIs and an uncertainty-aware training approach for classifying clinically significant PCa, to be applied in data-constrained settings such as local and small clinics. Method: Our approach involves a novel pipeline for translating unpaired 3.0T multi-parametric prostate MRIs to 1.5T, thereby augmenting the available training data. Additionally, we introduce an evidential deep learning approach to estimate model uncertainty and employ dataset filtering techniques during training. Furthermore, we propose a simple, yet efficient Evidential Focal Loss, combining focal loss with evidential uncertainty, to train our model effectively. Results: Our experiments demonstrate that the proposed method significantly improves the Area Under ROC Curve (AUC) by over 20% compared to the previous work (98.4% vs. 76.2%). Conclusions: Our proposed framework effectively translates and aligns public data with local data to increase the number of training data for deep models. Additionally, our proposed uncertainty estimation method enhances PCa detection performance. Providing prediction uncertainty to radiologists may aid in prioritizing uncertain cases, and expediting the diagnostic process effectively. Our code is available at https://github.com/med-i-lab/DT_UE_PCa.
keywords:
Deep Learning, Image Translation, Uncertainty Estimation, Prostate Cancer
*The majority of this work was completed while M.Zhou was at the Medical Informatics Laboratory, Queen’s University.
[…]
@@ 1 Introduction
Prostate Cancer (PCa) is a prevalent form of cancer among men [1], and the clinically significant PCa is defined by the Gleason score $>$ 6 or the histopathology ISUP grade $\geq$ 2 [2, 3]. The current PCa diagnosis procedure involves a combination of the prostate-specific antigen test and the histopathology analysis of the Transrectal Ultrasound-guided biopsy (TRUS) taken from 10-12 regions on the prostate gland [4, 5]. However, the histopathology analysis on TRUS can miss up to 20% of clinically significant PCa due to the limited number of biopsy samples [5, 1].
@@
Multi-parametric Magnetic Resonance Imaging (mp-MRI) has emerged as an effective alternative to TRUS for the early detection of PCa. mp-MRI uses a combination of anatomical and functional sequences of MRI that can further highlight the differences between normal and abnormal (cancer) cells. The evaluation and reporting guideline of prostate mp-MRI was first introduced in the Prostate Imaging Reporting and Data System (PI-RADS) [6, 7, 8]. The guideline provides a comprehensive scoring schema for suspicious prostate lesions and mp-MRI sequences. An extensive Prostate MRI imaging study (PROMIS) [9] reported that targeted biopsy using mp-MRI has higher sensitivity and negative-predictive value (NPV) but lower specificity compared to TRUS biopsy [9, 10, 11]. The study also showed that 27% of the patients did not need to undergo biopsy, had mp-MRI been used for screening. Although PROMIS provides strong practical implications for mp-MRI in PCa diagnosis, the low specificity indicates that mp-MRI can be plausibly improved by advanced analyses.
@@
In recent years, deep learning methods have emerged as a powerful tool for image classification tasks, and have provided promising performance in detecting and segmenting PCa on multi-parametric Prostate MRIs [12, 13, 14, 15, 16]. A more recent grand challenge, ProstateX [17], has further shown the ability of deep learning approaches in detecting clinically significant PCa on 3.0T mp-MRI data. Several groups have developed Convolutional Neural Network (CNN)-based models that achieve high performance for PCa classification [18, 19, 20, 17, 21, 5]. These methods have a great potential for clinical translations by highlighting abnormal lesions for radiologists during the PCa diagnostic process.
While deep learning has shown promising results in detecting PCa on mp-MRI, there are several substantial challenges in training and deploying deep models in clinics. Training deep models typically requires a large amount of data, which is not always available for local clinics with limited patient throughput. The alternative is to use pre-trained models or publicly available labeled data for training. However, the effectiveness of these models is significantly impacted by the differences in distribution between the public models/data and the images from local centers [5]. The primary contributing factor to this disparity is the field strength of MRI scanners.
While 3.0T scanner generally has higher image quality and spatial resolution [22] and most of the public datasets for prostate cancer are predominantly acquired with 3.0T scanners [23, 24, 25], 1.5T machines are considered standard of care and widely used in local clinical centers due to their cost-effectiveness [26]. Despite the preference for 3.0T MRI scanners, PI-RADS guidelines state that both 1.5T and 3.0T can provide adequate and reliable diagnostic examinations [27] and certain research indicates no statistically significant difference in clinical impact between 1.5T and 3.0T MRIs for diagnosis of prostate cancer [28, 29, 30]. Considering that over 80% of MRI systems in Canada operate at a 1.5T while exhibiting a higher rate of exams per hour than their 3.0T counterparts [31], 1.5T MRIs continue to be the standard in clinical care.
Therefore, to bridge the distribution gap between the public data and clinical application, a necessary step is to translate 3.0T MRI data to 1.5T. This translation process facilitates aligning the data distributions and allows for the training of classification models using both the translated data and local data. It is worth noting that there are ongoing efforts in the literature to address related challenges in federated learning [32, 33]. However, this aspect is not the focus of this study.
Besides, classical deep models are primarily designed for predicting labels during the inference of test data, irrespective of whether the test image is within or outside the training set distribution. These models lack the capability to discern data samples belonging to unrelated distributions [34] or to quantify their confidence in predictions. These limitations make models hard to interpret, raising concerns about the reliability of such models. Hence, reusing and deploying models for local PCa detection is challenging. Addressing the above limitations and drawbacks is crucial when integrating deep learning models into real clinical routines. Thus, two main questions arise in this context:
@@
    For small local clinical centers, can they take advantage of the extensive high-resolution 3.0T public MRI data to enhance the classification performance on their limited low-resolution local 1.5T MRI data?
@@
    When deploying models in clinical centers, can we provide supplementary information regarding the confidence of the model’s predictions, beyond the final result, to enhance the reliability and explainability of the models?
In this work, we aim to answer the two important questions outlined above. To this end, we propose a novel 2-stage learning framework for clinically significant PCa classification using multi-parametric, multi-center MRI data. Our framework not only aims to enhance classification performance but also provides an estimate of predictive confidence alongside the corresponding predicted label, introducing a valuable dimension to the model’s interpretability. In the first stage, we introduce a data preprocessing pipeline that translates prostate mp-MRI data from 3.0T to 1.5T via a Generative Adversarial Network (GAN) approach to increase the number of training samples. This step addresses the challenge of limited data in local clinics with low patient throughput (refer to Section 4.1). In the second stage, we propose an uncertainty-aware PCa classification approach. Specifically, we explore various model architectures to enhance classification performance, experiment with different strategies for combining multi-parametric MRI data, and leverage the co-teaching framework [35] to mitigate potential issues related to noisy labels (refer to Section 4.2.1). During the training phase, we incorporate dataset filtering using evidential uncertainty estimation [34] to eliminate training data samples with high prediction uncertainty, thereby enhancing the robustness of our models. Additionally, we extend the work of [34] by introducing a novel Evidential Focal Loss to optimize our classification models during training (see Section 4.2.2). The experimental results affirm the effectiveness of our proposed framework, demonstrating a significant improvement in classification performance compared to previous work. Our contributions not only advance the field of PCa detection but also underscore the potential of uncertainty-aware approaches in enhancing the reliability and interpretability of deep learning models in clinical settings.
@@
    We develop a GAN-based framework to translate unpaired prostate mp-MRIs from 3.0T to 1.5T, which we termed as domain transfer. This framework would align different data distributions and increase the number of training data for deep classification models. This is also the first attempt to translate prostate mp-MRIs in an unpaired manner.
@@
    We propose a novel loss function termed Evidential Focal Loss that can jointly compute the uncertainty for training samples and optimize the classification model. To the best of our knowledge, it is the first time that the original Focal Loss [36] is combined with the evidential uncertainty [34] for binary PCa classification.
@@
    By filtering the training samples based on their uncertainty value, our results outperform the state-of-the-art and improve the interpretability of model predictions. By providing confidence estimates for the predictions, radiologists can make informed decisions during the PCa diagnostic process and effectively expedite the process.
@@ 2.1 Domain Adaptation
Machine learning algorithms usually perform well when training and test data share the same distribution and feature space. However, in real-world applications, the distribution of test data often shifts, leading to biased or inaccurate predictions. In addition, it is time-consuming or infeasible to acquire new training data and fully repeat training steps. Domain Adaptation (DA) is an approach that addresses this issue by mitigating the dataset bias or domain shift problem caused by different distributions. There has been a lot of work on this topic in the past few years, which can be grouped into the following three general tasks [37]: (1) unsupervised DA tasks [38, 39, 40, 41, 42] focus on addressing the domain shift problem without requiring labeled target domain data; (2) semi-supervised DA tasks[43, 44, 45] aim to explore the partially labeled target domain data to further enhance the performance of domain adaptation algorithms; and (3) multi-source DA tasks [46, 47, 48] deal with scenarios where multiple source domains are available for adaptation. DA methods are often used to extract domain-invariant features for transferring knowledge between source and target domains. These methods incorporate various learning objectives with deep neural networks [49] for distribution matching: (1). Discrepancy Measurement-based methods aim to align feature distributions between two domains by fine-tuning deep models, e.g., using statistic criterion like Maximum Mean Discrepancy [50, 51, 52], and class criterion [53, 54, 55]. Some of these methods often require large labeled target domain data to diminish the domain shift problem, which is sometimes infeasible to get such medical data in the real-life scenario. (2). Adversarial-based methods aim to confuse domain discriminators from Generative Adversarial Networks (GANs) to enhance the invariant feature extraction [39, 56, 57]. One common scenario involves utilizing noise vectors, either with or without source images, to generate realistic target images while preserving the source features. However, training GANs are hard and sometimes results in generator degradation, e.g., mode collapse [58]. (3). Reconstruction-based methods, in addition to the general GANs approach from the above category, aim to reconstruct source-like images as an auxiliary task to preserve domain invariant features through an adversarial reconstruction paradigm [59, 60]. These methods usually have superior performance over the conventional GANs approach because they have an explicit reconstruction task to supervise the entire pipeline and make the training process more stable.
@@
CycleGAN [60] is one of the state-of-the-art unsupervised adversarial reconstruction-based methods that is widely used for unpaired image-to-image translation. Its cycle consistency loss ensures the pixel-level similarity between two images through a reconstruction task, i.e., the source image $s$ is translated to the target domain $\hat{s}$ and then translated back $\tilde{s}$, where it should be identical to the original image ($s=\tilde{s}$). However, a drawback of the cycle consistency loss lies in its stringent constraint on pixel-level similarity, which will degrade the performance of GANs in some certain tasks [61]. To address this limitation, the adversarial consistency loss GAN (ACL-GAN) [61] is proposed to replace the pixel-level similarity with the distance between distributions. This modification allows ACL-GAN to retain essential features from source images while overcoming the drawbacks associated with the strict cycle consistency constraint. Therefore, we adapt the ACL-GAN model and build our framework based on it.
@@
In medical imaging, domain shift problems usually fall into two variations: subject-related variation (age, gender, etc.), and acquisition-related variation (MRI vendor, field strength, imaging protocol, etc.) [62]. To solve such problem, one intuitive approach is to fine-tune a model that is pre-trained on the source domain with the new data from the target domain. [63] propose to use the pre-trained VGG model on the ImageNet dataset [64] to learn robust high-level features of natural images, and then fine-tune it on the labeled MR images for the Alzheimer’s Disease (AD) classification task to achieve state-of-the-art performance. Similarly, [65] study the impact of the fine-tuning techniques on the brain lesion segmentation task, demonstrating that fine-tuning with only a small number of target domain training samples can outperform models trained from scratch. Another approach is to use domain adaptation as an intermediate step to reduce variance in image acquisition parameters from both domains and then use it for downstream tasks. Researchers have attempted to address the problem of acquisition variation in MRI data for several years. [62] propose a feature-level representation learning method to either extract acquisition-invariant features or remove acquisition-variant features from paired 1.5T and 3.0T brain MRIs. The learned features are then used for a downstream classification task. However, obtaining paired 1.5T and 3.0T MRI data in real-life scenarios is impractical. Another way to align acquisition-invariant features is to synthesize images from different types of acquisition parameters using GAN-based adversarial reconstruction methods. GANs have been applied to perform cross-modality image translation between different medical images or generate synthetic images from random noise. The objective of such translation tasks is to retain the underlying structure while changing the appearance of the image [66]. Researchers have attempted to estimate images in the target modality from the source modality, such as MRI-CT translation [67, 68, 69, 70] and X-ray to CT translation [71, 72]. Other areas that have been explored include intra-modality translation, such as T1/T2-FLAIR translation [73, 74] and pure data augmentation by generating synthetic images from random noise vectors [75, 76, 77, 78]. However, most of the works do not consider the real clinical practicality, for example, T1/T2-FLAIR MRI translation may require paired training data, which is not feasible in real clinical settings. Generating synthetic images from noise does not take advantage of the publicly available data and ignores a-priori information. The current limitations provide great potential for unpaired image translation for medical images, which we employ in this work.